# From Birdwatch to Community Notes, from Twitter to X: four years of community-based content moderation

### v2

@@ RESULTS
The current design of Community Notes allows Contributors to classify a Post as either misleading or not misleading. The accompanying Notes serve to justify these classifications and provide additional context for the Post. Figure 2 (b) presents the number of Notes tagged as misleading or not misleading over the four-year period, revealing a consistent tendency for Contributors to label Posts as misleading more frequently.
Figure 2: Classification and the number of Notes per month. (a) The number of Notes and Ratings created within each month on a log scale. (b) The monthly number of Notes that classify the Posts as ”potentially misleading or misinformed” and ”not misleading” over four years on a log scale.
@@
A similar pattern emerges in the distribution of Ratings. Figure 3(b) shows the rank-plot of the number of Ratings submitted by each Contributor. Such patterns are ubiquitous in online open crowdsourcing platforms such as Wikipedia50, 51, citizen science projects52, Reddit53, and Urban Dictionary54, where a small number of ”power users” contribute a large proportion of the activities. This concentration of users can challenge the practical realisation of representation and openness in these systems.
@@
The Notes in our dataset were written on 1,016,673 distinct Posts. As shown in Figure 4(a), the distribution of Notes per Post is highly skewed: most Posts receive only a few Notes, while a small number attract substantial attention. The most annotated Post in the dataset has 90 Notes. This Post, made by Donald Trump on August 25, 2023, includes his mugshot alongside the caption: *”Election Interference, Never Surrender! DonaldjTrump.com”*. As of this writing, despite the large number of Notes on this Post, none of them have reached ”helpful” status, and therefore, no Note appears beneath it on the platform.
@@
We classified the 30 most frequently cited domains by political leaning using Ad Fontes Media’s 2023 ratings. Ad Fontes assigns each domain a score on a left–neutral–right spectrum (Left: –20 to –10, Neutral: –9 to +9, Right: +10 to +20)55. Figure 6 displays the top 30 cited domains and the number of times each was referenced. Crowd-sourced platforms such as Wikipedia, X, and YouTube rank among the most frequently cited. The bar colours represent the (U.S.-based) political leaning of each domain, revealing that most cited sources are classified as neutral, although several left-leaning domains also appear.
@@
Topic
Number
Number of Notes Topic Name Representative Words of Topic
@@
Figure 7: The size of the Ratings network over time. (a) The number of nodes for the helpful, unhelpful, and somewhat helpful networks for each month on the log scale. (b) The number of edges for the helpful, unhelpful, and somewhat helpful networks for each month on the log scale. Since the ”somewhat helpful” option was only introduced at the end of June 2021, the somewhat helpful network did not exist before that point.
Analysis of these rating networks offers valuable insight into the underlying community structure of the platform and its evolution over time (see, e.g., the work by Yasseri and Menczer (2023)9). As an illustrative example, Figure 8 presents visualisations of the three interaction networks—helpful, somewhat helpful, and unhelpful—constructed from ratings made in January 2023 as a representative month. This figure is intended primarily for illustration, and further analysis is required to fully interpret the structure of the networks. In particular, the unhelpful network (Figure 8(c)) warrants additional attention, since its edges represent negative interactions, yet they were treated as unsigned edges in the current visualisation.
Figure 8: Networks of January 2023. The network visualisation of (a) helpful ratings, (b) somewhat helpful ratings, and (c) unhelpful ratings made in January 2023. Node colours indicate cluster membership, with clusters identified using the Louvain algorithm56 with a Louvain resolution of 1. Only clusters containing more than 4% of all nodes are shown. In all networks, the size of each node is proportional to its weighted in-degree, representing the number of ratings a Contributor received during that month. The visualisations were created in Gephi $0.1$.
@@ DISCUSSION
The spread of misleading content on social media has had serious consequences, affecting public health outcomes57, 58, undermining electoral integrity59, 60, and even inciting violence61. In response, platforms have implemented a range of content moderation strategies, including third-party fact-checking and automated detection techniques. Introduced on X in January 2021, Community Notes is the first large-scale, community-driven content moderation system and has since begun gaining traction on other platforms as well62, 63.
Community Notes represents a fundamentally new approach to content moderation, distinguished not only by its crowdsourced nature but also by its epistemological foundation. As Augenstein et al. (2025) argue, there is a core philosophical distinction between Community Notes and traditional fact-checking systems43. Whereas traditional fact-checking is framed as the pursuit of an objective truth established through evidence, Community Notes are treated as subjective constructs reflecting personal opinion. This distinction has sparked debate over the functionality, credibility, and ethical implications of Community Notes as a moderation tool.
Community Notes leverages a broad pool of Contributors to evaluate a wide range of content and has shown promise in identifying misleading information14 and providing accurate, context-rich Notes13. Compared to traditional fact-checking interventions, Community Notes are less intrusive, giving users the autonomy to engage with or disregard them 43. Research also indicates that these Notes are perceived as more trustworthy than simple warning labels that lack explanation 22. Furthermore, the system’s algorithmic transparency has been shown to enhance user trust43.
@@
Fact-checking complex or high-stakes Posts often demands domain-specific expertise that most Contributors lack, and even knowledgeable Contributors may never encounter such Posts43. This limitation is compounded by the platform’s rating structure, which leaves a large share of Notes unpublished. Between January 23, 2021, and January 23, 2025, only 13.55% of Posts with at least one proposed Note ever received a “helpful” Note. Across the same period, 87.7% of all Notes remained in the “Needs More Ratings” category, and only 8.3% ultimately achieved “helpful” status and appeared beneath their respective Posts. Even then, the average delay was 26 hours—well past the point of peak visibility for most misleading Posts—greatly reducing the intervention’s practical impact64.
Several factors contribute to this bottleneck. First, ideological echo chambers limit exposure to diverse viewpoints, slowing the consensus-building needed for publication43. Second, political polarisation hinders agreement on contentious topics, making it harder for Notes to gain the cross-ideological support required to be deemed helpful13.
@@ METHODS
For Note content, Community Notes provides a single TSV file containing all Notes, with metadata including the Note ID, author ID, referenced Post ID, creation timestamp, the Contributor’s classification of the Post as misleading or not misleading, whether trustworthy sources were cited, and the full Note text65.
Ratings are available in multiple TSV files, each entry representing a single rating. These records contain the Note ID, the rater’s participant ID, perceived helpfulness, and reasons for the evaluation. The helpfulnessLevel field, introduced on June 30, 2021, replaced the earlier binary helpful/notHelpful label65.
@@
We used the pre-trained FastText language identification model66 to detect the language of each Note. English-language Notes were then processed through a custom text-cleaning pipeline optimised for topic modelling. This pipeline included HTML decoding, normalisation of accented characters, expansion of contractions, and removal of URLs, user mentions, hashtags, and non-alphabetic characters. We then removed stopwords, filtered out short tokens, and applied lemmatisation. Notes reduced to empty text after preprocessing were excluded to ensure clean input for embedding generation and topic modelling.
@@
We applied BERTopic67 for topic modelling. BERTopic is a modular framework comprising sequential stages, each customisable with different methodological components. The process begins with generating semantic embeddings of the text, followed by dimensionality reduction and clustering of these representations to identify distinct topics. Each cluster is then linked to its most representative terms to aid interpretability. In an optional post-processing step, semantically similar topics can be merged to produce a more concise and coherent final set of topics.
We used the pre-trained all-MiniLM-L6-v2 sentence transformer68, 69 to embed the preprocessed Community Notes text into a 384-dimensional vector space. To handle the large dataset size, we first applied Principal Component Analysis (PCA)70 for dimensionality reduction (see Table S4), followed by Uniform Manifold Approximation and Projection (UMAP)71. The UMAP initialisation parameters are listed in Table S5.
Clustering was conducted using Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN)72, which automatically estimates the number of clusters, making it well-suited for topic modelling where the number of topics is not known a priori. HDBSCAN is robust to noise and integrates effectively with UMAP-reduced data. Key hyperparameters for HDBSCAN are detailed in Table S6. We excluded a noise cluster from the list of topics. For the overall model parameters, see Table S7.
@@
To examine Contributor interaction dynamics within Community Notes, we constructed monthly networks based on the Ratings.
For the study period, January 2021 to January 2025, we created 49 monthly rating files, each used to generate a corresponding interaction network.
Each network is represented as a directed graph, where nodes correspond to Community Notes Contributors. A directed edge from a rater to a Note writer indicates that the rater evaluated the writer’s Note. Each edge contains three attributes—helpful, unhelpful, and somewhat helpful—representing the number of each respective rating the rater assigned to that Note writer during the month. These complete monthly graphs, referred to as the ”whole networks”, are available for download.
@@
This study did not generate new unique reagents.
@@ ACKNOWLEDGMENTS
Literature review, S.M., K.S., H.C., M.D., S.G., and T.Y.; data analysis, S.M., N.C., H.D., K.S.; drafting the paper, S.M., N.C., H.D., T.Y., supervision, T.Y; All authors approved the manuscript.
@@
During the preparation of this work, the authors used ChatGPT 4.0 to improve the writing style of this article. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
SUPPLEMENTAL INFORMATION INDEX
Document S1. Tables S1-S7 and their captions in the PDF.
Document S2. Table S8. Full list of topics and their representative words sorted by topic size is available at  https://doi.org/10.5281/zenodo.16761304. .
SUPPLEMENTAL INFORMATION

### v3

@@ RESULTS
The current design of Community Notes allows Contributors to classify a Post as either misleading or not misleading. The accompanying Notes serve to justify these classifications and provide additional context for the Post. Figure 2 (b) presents the number of Notes tagged as misleading or not misleading over the four-year period.
Figure 2: Classification and the number of Notes per month. (a) The number of Notes and Ratings created within each month on a log scale. (b) The monthly number of Notes that classify the Posts as “potentially misleading or misinformed” and “not misleading” over four years on a log scale.
@@
A similar pattern emerges in the distribution of Ratings. Figure 3(b) shows the rank-plot of the number of Ratings submitted by each Contributor. Such patterns are ubiquitous in online open crowdsourcing platforms such as Wikipedia (Ortega et al. (2008); Yasseri and Kertész (2013)), citizen science projects (Sauermann and Franzoni (2015)), Reddit (Glenski et al. (2017)), and Urban Dictionary (Nguyen et al. (2018)), where a small number of “power users” contribute a large proportion of the activities. This concentration of users can challenge the practical realisation of representation and openness in these systems.
@@
The Notes in our dataset were written on 1,016,673 distinct Posts. As shown in Figure 4(a), the distribution of Notes per Post is highly skewed: most Posts receive only a few Notes, while a small number attract substantial attention. The most annotated Post in the dataset has 90 Notes. This Post, made by Donald Trump on August 25, 2023, includes his mugshot alongside the caption: “Election Interference, Never Surrender! DonaldjTrump.com”. As of this writing, despite the large number of Notes on this Post, none of them have reached “helpful” status, and therefore, no Note appears beneath it on the platform.
@@
We classified the 30 most frequently cited domains by political leaning using Ad Fontes Media’s 2023 ratings. Ad Fontes assigns each news source a score along a left–neutral–right spectrum (Left: –42 to –10, Neutral: –9 to +9, Right: +10 to +42). These ratings are determined by panels of trained analysts with diverse political ideologies who evaluate samples of articles from each source and manually assign bias scores based on language, framing, and political positioning (Ad Fontes Media, Inc. (2025)).
Figure 6 displays the top 30 cited domains and the number of times each was referenced. Crowd-sourced platforms such as Wikipedia, X, and YouTube rank among the most frequently cited. The bar colours represent the (U.S.-based) political leaning of each domain, revealing that most cited sources are classified as neutral, although several left-leaning domains also appear.
@@
Topic Number Number of Notes Topic Name Representative Words of Topic
@@
Figure 7: The size of the Ratings network over time. (a) The number of nodes for the helpful, unhelpful, and somewhat helpful networks for each month on the log scale. (b) The number of edges for the helpful, unhelpful, and somewhat helpful networks for each month on the log scale. Since the “somewhat helpful” option was only introduced at the end of June 2021, the somewhat helpful network did not exist before that point.
Analysis of these rating networks offers valuable insight into the underlying community structure of the platform and its evolution over time (see, e.g., the work by Yasseri and Menczer (2023)). As an illustrative example, Figure 8 presents visualisations of the three interaction networks—helpful, somewhat helpful, and unhelpful—constructed from ratings made in January 2023 as a representative month. This figure is intended primarily for illustration, and further analysis is required to fully interpret the structure of the networks. In particular, the unhelpful network (Figure 8(c)) warrants additional attention, since its edges represent negative interactions, yet they were treated as unsigned edges in the current visualisation.
Figure 8: Networks of January 2023. The network visualisation of (a) helpful ratings, (b) somewhat helpful ratings, and (c) unhelpful ratings made in January 2023. Node colours indicate cluster membership, with clusters identified using the Louvain algorithm (Blondel et al. (2008)) with a Louvain resolution of 1. Only clusters containing more than 4% of all nodes are shown. In all networks, the size of each node is proportional to its weighted in-degree, representing the number of ratings a Contributor received during that month. The visualisations were created in Gephi $0.1$.
@@ DISCUSSION
The spread of misleading content on social media has had serious consequences, affecting public health outcomes (Loomba et al. (2021); Enders et al. (2022)), undermining electoral integrity (Allcott and Gentzkow (2017); Greene et al. (2021)), and even inciting violence (Klein (2025)). In response, platforms have implemented a range of content moderation strategies, including third-party fact-checking and automated detection techniques. Introduced on X in January 2021, Community Notes is the first large-scale, community-driven content moderation system and has since begun gaining traction on other platforms as well (Kaplan (2025); TikTok Newsroom (2025)).
Community Notes represents a fundamentally new approach to content moderation, distinguished not only by its crowdsourced nature but also by its epistemological foundation. As Augenstein et al. (2025) argue, there is a core philosophical distinction between Community Notes and traditional fact-checking systems. Whereas traditional fact-checking is framed as the pursuit of an objective truth established through evidence, Community Notes are treated as subjective constructs reflecting personal opinion. This distinction has sparked debate over the functionality, credibility, and ethical implications of Community Notes as a moderation tool.
Community Notes leverages a broad pool of Contributors to evaluate a wide range of content and has shown promise in identifying misleading information (Martel et al. (2024)) and providing accurate, context-rich Notes (Saeed et al. (2022)). Compared to traditional fact-checking interventions, Community Notes are less intrusive, giving users the autonomy to engage with or disregard them (Augenstein et al. (2025)). Research also indicates that these Notes are perceived as more trustworthy than simple warning labels that lack explanation (Drolsbach et al. (2024)). Furthermore, the system’s algorithmic transparency has been shown to enhance user trust (Augenstein et al. (2025)).
@@
Fact-checking complex or high-stakes Posts often demands domain-specific expertise that most Contributors lack, and even knowledgeable Contributors may never encounter such Posts (Augenstein et al. (2025)). This limitation is compounded by the platform’s rating structure, which leaves a large share of Notes unpublished. Between January 23, 2021, and January 23, 2025, only 13.55% of Posts with at least one proposed Note ever received a “helpful” Note. Across the same period, 87.7% of all Notes remained in the “Needs More Ratings” category, and only 8.3% ultimately achieved “helpful” status and appeared beneath their respective Posts. Even then, the average delay was 26 hours—well past the point of peak visibility for most misleading Posts—greatly reducing the intervention’s practical impact (Truong et al. (2025a)).
Several factors contribute to this bottleneck. First, ideological echo chambers limit exposure to diverse viewpoints, slowing the consensus-building needed for publication (Augenstein et al. (2025)). Second, political polarisation hinders agreement on contentious topics, making it harder for Notes to gain the cross-ideological support required to be deemed helpful (Saeed et al. (2022)).
@@ METHODS
For Note content, Community Notes provides a single TSV file containing all Notes, with metadata including the Note ID, author ID, referenced Post ID, creation timestamp, the Contributor’s classification of the Post as misleading or not misleading, whether trustworthy sources were cited, and the full Note text (X Documentation (2025a)).
Ratings are available in multiple TSV files, each entry representing a single rating. These records contain the Note ID, the rater’s participant ID, perceived helpfulness, and reasons for the evaluation. The helpfulnessLevel field, introduced on June 30, 2021, replaced the earlier binary helpful/notHelpful label (X Documentation (2025a)).
@@
We used the pre-trained FastText language identification model (Joulin et al. (2016)) to detect the language of each Note. English-language Notes were then processed through a custom text-cleaning pipeline optimised for topic modelling. This pipeline included HTML decoding, normalisation of accented characters, expansion of contractions, and removal of URLs, user mentions, hashtags, and non-alphabetic characters. We then removed stopwords, filtered out short tokens, and applied lemmatisation. Notes reduced to empty text after preprocessing were excluded to ensure clean input for embedding generation and topic modelling.
@@
We applied BERTopic (Grootendorst (2022)) for topic modelling. BERTopic is a modular framework comprising sequential stages, each customisable with different methodological components. The process begins with generating semantic embeddings of the text, followed by dimensionality reduction and clustering of these representations to identify distinct topics. Each cluster is then linked to its most representative terms to aid interpretability. In an optional post-processing step, semantically similar topics can be merged to produce a more concise and coherent final set of topics.
We used the pre-trained all-MiniLM-L6-v2 sentence transformer (Reimers and Gurevych (2019); Wang et al. (2020)) to embed the preprocessed Community Notes text into a 384-dimensional vector space. To handle the large dataset size, we first applied Principal Component Analysis (PCA) (Pedregosa et al. (2011)) for dimensionality reduction (see Table S4), followed by Uniform Manifold Approximation and Projection (UMAP) (McInnes et al. (2018)). The UMAP initialisation parameters are listed in Table S5.
Clustering was conducted using Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) (McInnes et al. (2017)), which automatically estimates the number of clusters, making it well-suited for topic modelling where the number of topics is not known a priori. HDBSCAN is robust to noise and integrates effectively with UMAP-reduced data. Key hyperparameters for HDBSCAN are detailed in Table S6. We excluded a noise cluster from the list of topics. For the overall model parameters, see Table S7.
@@
To examine Contributor interaction dynamics within Community Notes, we constructed monthly networks based on the Ratings. For the study period, January 2021 to January 2025, we created 49 monthly rating files, each used to generate a corresponding interaction network. Each network is represented as a directed graph, where nodes correspond to Community Notes Contributors. A directed edge from a rater to a Note writer indicates that the rater evaluated the writer’s Note. Each edge contains three attributes—helpful, unhelpful, and somewhat helpful—representing the number of each respective rating the rater assigned to that Note writer during the month. These complete monthly graphs, referred to as the “whole networks”, are available for download.
@@
This study did not generate new, unique reagents.
@@ ACKNOWLEDGMENTS
The authors would like to thank all the researchers whose work was reviewed in this paper for their valuable comments and feedback, particularly Nicolas Pröllochs.
@@
Introduction, S.M. and T.Y.; Literature review, S.M., H.C., M.D., S.G., and N.C; Data merging, K.S.; Network extraction, S.M., K.S., and M.D.; Overall statistics, S.M., and H.C.; Language statistics, N.C., and H.D.; Source analysis, N.C., and H.D.; Topic modelling, N.C., and H.D.; Network visualisation, S.M.; Discussion, S.M., and T.Y.; Data disposition, S.M.; Supervision, coordination and conception, T.Y. and S.M.; All authors approved the manuscript.
@@
During the preparation of this work, the authors used ChatGPT 4.1 to improve the writing style of this article. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
SUPPLEMENTAL INFORMATION for
From Birdwatch to Community Notes, from Twitter to X:
four years of community-based content moderation
Saeedeh Mohammadi, Narges Chinichian, Hannah Doyal, Kristina Skutilova, Hao Cui, Michele d’Errico, Siobhan Grayson, Taha Yasseri
@@
Table S8: Full list of topics and their representative words sorted by topic size in an .xls file.